An attacker at the base of a castle wall 3.70 m hi
Answer #1

by the energy conservrtion

height it travell is 3.7-1.6=2.1m

K.E initial + P.Einitial = K.Efinal + P.E.Final

0.5mvi2 +mgh1 = 0.5mvf2 + mgh2

m will be cancell from all over the eqaution

0.5*7.52 + 9.8*1.6 = 0.5vf2 + 9.8*3.7

Vf = 3.88458 m/s

yes it reach the top of casstle

part b

speed at top

vf =3.8845 m/s

change in speed=7.5-3.8845=3.6155 m/s

part c

now energy equation

by cancelling m

0.5vtop2 +ghtop = 0.5vground2 + ghlow

0.5*7.52 + 9.8*3.7 = 0.5vground2 + 9.8*1.6

vground =9.8695 m/s

change in speed =9.8695-7.5=2.369 m/s

part d

so it is not same